OpenStudy (anonymous):

A golf ball is hit from the ground with an initial velocity of 208 feet per second. Assume the starting height of the ball is 0 feet. How long will it take the golf ball to hit the ground? Using H(t)=-16t^2+vt+s

OpenStudy (anonymous):

s=0 since initial height =0; velocity =208 f/sec while hitting the grnd 0=-16t^2+208t t=208/16=13 sec